OpenStudy (anonymous):

Can anyone help me with Calculus? A rescue boat uses a long towline and a winch located twelve feet off the face of the water to receive a small yacht in distress. on the rescue boat the towline is wrapped once around a 3 - foot radius cylinder that retrieves the line by turning at a rate of 2 revolutions per minute. Determine the rate at which the yacht is approaching the tugboat when it is 50 ft away .

OpenStudy (anonymous):

Would I use the formula for volume of a cylinder or use Pythagorean theorem?

OpenStudy (anonymous):

pythagorean theorem

OpenStudy (anonymous):

the cylindrical shape is irrelevant

OpenStudy (anonymous):



OpenStudy (anonymous):

b = 12

OpenStudy (anonymous):

dy/dt is 3-foot?

OpenStudy (anonymous):

3 is not important

OpenStudy (anonymous):

a = 50

OpenStudy (anonymous):

c = sqrt(a^2+b^2)

OpenStudy (anonymous):

you want to find da/dt

OpenStudy (anonymous):

is there a constant?

OpenStudy (anonymous):

c = sqrt(a^2+b^2) -> c^2 = a^2 + b^2 -> c^2 - b^2 = a^2

OpenStudy (anonymous):

*derivative* 2c(dc/dt) - 2b(db/dt) = 2a(da/dt)

OpenStudy (anonymous):

db/dt = 0, so 2c(dc/dt) - 2b(0) = 2a(da/dt) -->2c(dc/dt) = 2a(da/dt) -> c(dc/dt) = a(da/dt)

OpenStudy (anonymous):

is it sqrt 2644?

OpenStudy (anonymous):

2c(dc/dt) = 2a(da/dt) -> (da/dt) = c/a *(dc/dt)

OpenStudy (anonymous):



OpenStudy (anonymous):

hypotenuse is sqrt2644

OpenStudy (anonymous):

just trig

OpenStudy (anonymous):

The ? determines the length of the tow line to the yacht. You will need to determine the circumference of the cylinder as that will give yu the number of feet of tow line that is being pulled in each minute.

OpenStudy (anonymous):

how do i know which one is dx/dt...

OpenStudy (anonymous):

(da/dt) = c/a *(dc/dt) This is where the cylinder comes in; for dc/dt we know 2 revolutions per min, which means total length of rope pulled is 2*[2pi(3)] = dc/dt

OpenStudy (anonymous):

x is a

OpenStudy (anonymous):

dc/dt = 12pi

OpenStudy (anonymous):

da/dt = c/a *(dc/dt) = sqrt2644 / 50 * 12pi

OpenStudy (anonymous):

da/dt = dx/dt

OpenStudy (anonymous):

x=a y=b

OpenStudy (anonymous):

TLDR; dx/dt = sqrt2644 / 50 * 12pi

OpenStudy (anonymous):

how would I know da/dt=c/a*dc/dt

OpenStudy (anonymous):

from differentiating the pythagorean theorem with respect to t

OpenStudy (anonymous):

(da/dt) = c/a *(dc/dt) This is where the cylinder comes in; for dc/dt we know 2 revolutions per min, which means total length of rope pulled is 2*[2pi(3)] = dc/dt how did u get that from 2revolutiongs per min?